# Double Robustness of Local Projections

### v2

@@ 4.3 Bias-aware inference
where the bias-aware critical value $\operatorname{cv}_{1-a}(b)$ is given by the number such that $r(b;\operatorname{cv}_{1-a}(b))=a$, and $r(\cdot,\cdot)$ is defined in Corollary 3.1. By construction, this bias-aware confidence interval has correct (but potentially conservative) asymptotic coverage.
@@
Impose Assumption 2.1, $\zeta=1/2$, and $\operatorname{aVar}(\hat{\delta}_{h})>0$. Then
@@
The result follows immediately from Propositions 3.2 and 4.1. ∎
It turns out, however, that a very tight bound $M$ on the signal-to-noise ratio is required for the bias-aware VAR interval to be shorter than the LP interval. Figure 4.3 plots the relative interval length as a function of the relative asymptotic standard deviation of VAR and LP, for a significance level of $a=10\%$ and for different misspecification bounds $M$. The figure shows that $M$ has to be quite small—apparently below 1—for the bias-aware VAR length to dominate the LP length regardless of the DGP and horizon. Even for $M=1.5$, bias-aware VAR is at best only moderately shorter than LP. Finally, for values of $M$ above 2 (corresponding to a noise-to-signal ratio above $4/T$), bias-aware VAR is dominated by LP.
Figure 4.3: Relative length of bias-aware VAR confidence interval vs. conventional LP interval. Significance level $a=10\%$. Horizontal axis: relative asymptotic standard deviation of LP vs. VAR. Different lines: different bounds $M$ on $\|\alpha(L)\|$. The solid horizontal line marks the value 1.
In Section A.2 we furthermore show that the conventional LP confidence interval is at worst slightly wider than a more efficient bias-aware confidence interval centered at the model averaging estimator $\hat{\theta}_{h}(\omega)=\omega\hat{\beta}_{h}+(1-\omega)\hat{\delta}_{h}$, introduced in Corollary 4.2 above. Even if the weight $\omega$ is chosen to optimize confidence interval length, the gains relative to the LP interval are very small when $M\geq 2$ (corresponding to a noise-to-signal ratio above $4/T$).
@@ 4.4 Inference on multiple impulse responses
Since the least favorable MA polynomial derived in Section 4.2 depends on the horizon $h$ of interest, one might hope that VARs would not be as prone to bias and thereby undercoverage if interest centers on multiple impulse responses. Unfortunately, we now show that this is not the case by generalizing the worst-case bias formula to the multi-dimensional case and deriving the worst-case coverage of the Wald confidence ellipsoid.
@@
We now consider inference on any combination of impulse responses for various horizons $h$, response variables $i^{*}$, and shocks $j^{*}$. When referring to impulse responses and estimators of these, we need to make the response variable and shock explicit in the notation. Thus, we write $\theta_{i^{*},j^{*},h,T}$, $\hat{\beta}_{i^{*},j^{*},h}$, and $\hat{\delta}_{i^{*},j^{*},h}$, with the definitions being the same as in Section 2. Let $k$ denote the total number of impulse responses of interest. We refer to the list of impulse responses by the collection of triples $\{(i_{a}^{*},j_{a}^{*},h_{a})\}_{a=1}^{k}$ indexing the response variable, shock variable, and horizon, respectively. Define the $k$-dimensional vectors of true impulse responses and LP and VAR estimators:
$\boldsymbol{\theta}_{T}\equiv\begin{pmatrix}\theta_{i_{1}^{*},j_{1}^{*},h_{1},T}\\ \vdots\\ \theta_{i_{k}^{*},j_{k}^{*},h_{k},T}\end{pmatrix},\quad\hat{\boldsymbol{\beta}}\equiv\begin{pmatrix}\hat{\beta}_{i_{1}^{*},j_{1}^{*},h_{1}}\\ \vdots\\ \hat{\beta}_{i_{k}^{*},j_{k}^{*},h_{k}}\end{pmatrix},\quad\hat{\boldsymbol{\delta}}\equiv\begin{pmatrix}\hat{\delta}_{i_{1}^{*},j_{1}^{*},h_{1}}\\ \vdots\\ \hat{\delta}_{i_{k}^{*},j_{k}^{*},h_{k}}\end{pmatrix}.$
It follows from Propositions 3.1 and 3.2 that
$\sqrt{T}\begin{pmatrix}\hat{\boldsymbol{\beta}}-\boldsymbol{\theta}_{T}\\ \hat{\boldsymbol{\delta}}-\boldsymbol{\theta}_{T}\end{pmatrix}\stackrel{{\scriptstyle d}}{{\to}}N\left(\begin{pmatrix}0_{k\times 1}\\ \operatorname{aBias}(\hat{\boldsymbol{\delta}})\end{pmatrix},\begin{pmatrix}\operatorname{aVar}(\hat{\boldsymbol{\beta}})&\operatorname{aCov}(\hat{\boldsymbol{\beta}},\hat{\boldsymbol{\delta}})\\ \operatorname{aCov}(\hat{\boldsymbol{\delta}},\hat{\boldsymbol{\beta}})&\operatorname{aVar}(\hat{\boldsymbol{\delta}})\end{pmatrix}\right),$ (4.2)
for a $k$-dimensional vector $\operatorname{aBias}(\hat{\boldsymbol{\delta}})$ (defined in the proof of Proposition 4.2 below) and $k\times k$ matrices $\operatorname{aVar}(\hat{\boldsymbol{\beta}})$, $\operatorname{aVar}(\hat{\boldsymbol{\delta}})$, and $\operatorname{aCov}(\hat{\boldsymbol{\beta}},\hat{\boldsymbol{\delta}})$ given in Corollary A.2 in Section A.3. This corollary also implies that the difference $\hat{\boldsymbol{\beta}}-\hat{\boldsymbol{\delta}}$ is asymptotically independent of $\hat{\boldsymbol{\delta}}$, which is not surprising given the general arguments of Hausman (1978) and the facts that (i) the asymptotic variances of the estimators are the same as in the model with $\alpha(L)=0$ and (ii) the VAR estimator is the quasi-MLE in such a model. It follows that $\operatorname{aVar}(\hat{\boldsymbol{\beta}})\geq\operatorname{aVar}(\hat{\boldsymbol{\delta}})$ in the positive semidefinite sense.
Worst-case bias.
The following result generalizes the univariate worst-case bias formula in Proposition 4.1.
Proposition 4.2.
Impose Assumption 2.1, with part (iii) holding for all shock indices $j_{1}^{*},\dots,j_{k}^{*}$, and let $\zeta=1/2$. Let $R$ be a constant matrix with $k$ rows. Then
$\max_{\alpha(L)\colon\|\alpha(L)\|\leq M}\|R\operatorname{aBias}(\hat{\boldsymbol{\delta}})\|^{2}=M^{2}\lambda_{\max}\!\left(R[\operatorname{aVar}(\hat{\boldsymbol{\beta}})-\operatorname{aVar}(\hat{\boldsymbol{\delta}})]R^{\prime}\right),$
where $\lambda_{\max}(B)$ denotes the largest eigenvalue of the matrix $B$.
Proof.
See Section B.3. ∎
The proposition shows that the worst-case squared norm of the bias of the VAR estimator $R\hat{\boldsymbol{\delta}}$ of $R\boldsymbol{\theta}_{T}$ is a function of two simple quantities: the bound $M$ on misspecification, and the largest eigenvalue of the difference $\operatorname{aVar}(R\hat{\boldsymbol{\beta}})-\operatorname{aVar}(R\hat{\boldsymbol{\delta}})$ between the variance-covariance matrices for the LP and VAR estimators. The latter eigenvalue equals $\max_{\|\varsigma\|=1}\{\operatorname{aVar}(\varsigma^{\prime}R\hat{\boldsymbol{\beta}})-\operatorname{aVar}(\varsigma^{\prime}R\hat{\boldsymbol{\delta}})\}$, i.e., the largest efficiency gain for VAR over LP across all linear combinations (with norm 1) of the estimated parameters. We thus see that there is still no free lunch: the worst-case bias is non-negligible if the VAR offers efficiency gains for any linear combination of the parameters of interest, echoing our univariate results. When $R$ is a row vector, then the proposition implies that our conclusions from Section 4.2 extend to inference on any linear combination of impulse responses (across horizons, variables, and/or shocks). In particular, the VAR confidence interval has fragile coverage even if the target parameter is the integral (i.e., sum) of impulse responses across multiple horizons, as in the fiscal multiplier applications reviewed in Ramey (2016).
Worst-case coverage of confidence ellipsoid.
We next derive the coverage of the conventional Wald confidence ellipsoid based on the VAR estimator. The level-($1-a$) confidence ellipsoid is given by
$\operatorname{CE}(\hat{\boldsymbol{\delta}})\equiv\left\{\tilde{\boldsymbol{\theta}}\in\mathbb{R}^{k}\colon T(\hat{\boldsymbol{\delta}}-\tilde{\boldsymbol{\theta}})^{\prime}\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1}(\hat{\boldsymbol{\delta}}-\tilde{\boldsymbol{\theta}})\leq\chi_{1-a,k}^{2}\right\},$
where $\chi_{1-a,k}^{2}$ is the $1-a$ quantile of the $\chi^{2}$ distribution with $k$ degrees of freedom.
Corollary 4.6.
Impose Assumption 2.1, with part (iii) holding for all shock indices $j_{1}^{*},\dots,j_{k}^{*}$, and let $\zeta=1/2$. Assume also that $\operatorname{aVar}(\hat{\boldsymbol{\delta}})$ is non-singular. Then
$\min_{\alpha(L)\colon\|\alpha(L)\|\leq M}\lim_{T\to\infty}P(\boldsymbol{\theta}_{T}\in\operatorname{CE}(\hat{\boldsymbol{\delta}}))=F_{k}\left(\chi_{1-a,k}^{2};M^{2}\left[\lambda_{\max}(\operatorname{aVar}(\hat{\boldsymbol{\beta}})\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1})-1\right]\right),$
where $F_{k}(x;c)$ is the cumulative distribution function, evaluated at point $x\geq 0$, of a non-central $\chi^{2}$ distribution with $k$ degrees of freedom and non-centrality parameter $c\geq 0$.
Proof.
See Section B.8. ∎
The worst-case coverage probability of the VAR confidence ellipsoid depends on three scalars: the bound $M$ on misspecification, the dimension $k$ of the ellipsoid, and the “multivariate relative standard error”
$\sqrt{\lambda_{\min}(\operatorname{aVar}(\hat{\boldsymbol{\delta}})\operatorname{aVar}(\hat{\boldsymbol{\beta}})^{-1})}=[\lambda_{\max}(\operatorname{aVar}(\hat{\boldsymbol{\beta}})\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1})]^{-1/2}=\min_{\varsigma\in\mathbb{R}^{k}}\sqrt{\operatorname{aVar}(\varsigma^{\prime}\hat{\boldsymbol{\delta}})/\operatorname{aVar}(\varsigma^{\prime}\hat{\boldsymbol{\beta}})}.$
Again, the worst-case coverage distortion is an increasing function of the largest efficiency gain for VAR over LP across all linear combinations of the impulse responses. Since VAR impulse response estimates are often highly correlated across horizons, this suggests that the VAR undercoverage can in fact be particularly severe in the multivariate case.
Figures 4.4 and 4.5 show that the worst-case coverage probability of the confidence ellipsoid can be very poor even for small amounts of misspecification. The panels in the figures correspond to choices of the dimension $k\in\{2,5,10\}$ or the bound $M\in\{1,1.5,2\}$ on misspecification, respectively. Evidently, the coverage distortions can be severe for all of the $k$ considered in these plots even when $M=1$ (corresponding to a noise-to-signal ratio of $1/T$), especially if the multivariate relative standard error $\sqrt{\lambda_{\min}(\operatorname{aVar}(\hat{\boldsymbol{\delta}})\operatorname{aVar}(\hat{\boldsymbol{\beta}})^{-1})}\leq 1/4$. It further appears that the worst-case coverage distortion is monotonically decreasing in $k$, holding everything else constant; however, loosely speaking, the larger is $k$, the larger is the “chance” that there is some linear combination of the parameters for which VAR is much more efficient than LP, instead yielding a larger coverage distortion. The multivariate case thus overall echoes and reinforces our conclusions from the earlier univariate analysis.
Figure 4.4: Worst-case asymptotic coverage probability of the conventional 90% VAR confidence ellipsoid. Different panels: choices of the number $k$ of parameters in the ellipsoid. Horizontal axes: square root of smallest eigenvalue of the ratio of the VAR and LP variance-covariance matrices. Different lines: different bounds $M$ on $\|\alpha(L)\|$. The solid horizontal lines mark the nominal coverage probability $1-a=90\%$.
Figure 4.5: Worst-case asymptotic coverage probability of the conventional 90% VAR confidence ellipsoid. Different panels: different bounds $M$ on $\|\alpha(L)\|$. Horizontal axes: square root of smallest eigenvalue of the ratio of the VAR and LP variance-covariance matrices. Different lines: choices of the number $k$ of parameters in the ellipsoid. The solid horizontal lines mark the nominal coverage probability $1-a=90\%$.
5 Simulations
We now show that our asymptotic results accurately reflect the finite-sample properties of LP and VAR procedures. To do so, we consider a standard structural macroeconomic model as our DGP; in LABEL:app:simul_arma we present further illustrative results from a simple univariate model.
Framework.
Our DGP is the well-known structural macroeconomic model of Smets and Wouters (2007). This environment is ideal for our purposes, as it allows us to closely mimic applied macroeconometric practice: the model is rich enough to match quite well the second-moment properties of standard macroeconomic time series data, yet at the same time features the interpretable, structural shocks typically studied in applied work. We solve the model at the posterior mode estimated by Smets and Wouters. For our main exercise we will assume that the econometrician observes the wage cost-push shock, inflation, wages, and total hours worked. The impulse response function of interest is that of inflation with respect to the cost-push shock. In addition to being topical (e.g., see the recent work of Bernanke and Blanchard, 2023), this exercise is well-suited to cleanly illustrate the potential bite of our theoretical conclusions: wage cost-push shocks in the model of Smets and Wouters follow an ARMA($1$,$1$) process, and so short-lag VARs may be subject to material biases. Results for a monetary shock specification are reported in LABEL:app:sim_observedmp.
@@ Results.
$p$ $M$ $\frac{M^{2}}{1+M^{2}}$
1 34.053 0.999
2 5.412 0.967
4 3.225 0.912
8 1.893 0.782
12 1.324 0.637
20 0.915 0.456
40 0.614 0.274
Table 5.1: $M$ and $\frac{M^{2}}{1+M^{2}}$ as functions of $p$ in the model of Smets and Wouters, with the researcher observing the cost-push shock, inflation, wages, and hours worked, and estimating a VAR($p$).
@@
Lag length $p=4$
Lag length $p=8$
Figure 5.1: Coverage probabilities (left) and median confidence interval length (right) for VAR (red) and LP (blue) confidence intervals computed via the delta method or bootstrap (the latter are indicated with subscript “b” in the figure legends). DGP: Smets and Wouters (2007), cost-push shock. Lag length $p$ is selected using the AIC for the top panel, and fixed at $p=4$ and $p=8$ for the middle and bottom panels, respectively.
We begin by quantifying the amount of VAR misspecification in this DGP. Given a choice of lag length $p$, we represent the Smets and Wouters model in VARMA($p,\infty$) form (2.1) with $T=240$ and $\zeta=1/2$, where the VAR coefficients are selected to minimize the population sum of squared residuals.888 Let $y_{t}=C(L)u_{t}$ be the Wold decomposition of the observed data. Denote the linear least-squares projection coefficients of $y_{t}$ onto $p$ of its lags as $A_{\ell}(p)$, $\ell=1,2,\dots,p$, and define $A(L;p)\equiv I-\sum_{\ell=1}^{p}A_{\ell}(p)$. Letting $\tilde{C}(L;p)=\sum_{\ell=0}^{\infty}\tilde{C}_{\ell}(p)L^{\ell}\equiv A(L;p)C(L)$, we obtain a representation $y_{t}=\sum_{\ell=1}^{p}A_{\ell}(p)y_{t-\ell}+\sum_{\ell=0}^{\infty}\tilde{C}_{\ell}(p)u_{t-\ell}$. It is straightforward to orthogonalize the innovations in this representation and write it in the form (2.1), thus yielding the implied MA polynomial $\alpha(L)$. Table 5.1 shows, as a function of $p$, the total degree of misspecification $M=\|\alpha(L)\|$ as well as $M^{2}/(1+M^{2})$, the minimax optimal ex ante model averaging weight on LP in Corollary 4.2. As anticipated, the larger $p$, the smaller $M$. Importantly, however, $M$ only declines extremely slowly with the lag length $p$. In particular, for lag lengths typical in applied practice (for quarterly data), misspecification is material, with $M\approx 3.23$ for a standard lag length of $p=4$, and $M\approx 1.89$ even for a long lag length of $p=8$. This suggests that—depending on the shape of $\alpha(L)$—there is potential for misspecification to substantially affect VAR inference.
[…]
LABEL:app:sim_observedmp shows that these conclusions extend to monetary shock specifications. We furthermore show that the VAR coverage distortions in the actual Smets and Wouters DGP are comparable to those in a slightly perturbed DGP where we replace the model-implied MA lag polynomial $\alpha(L)$ by the theoretical least favorable one $\alpha^{\dagger}(L)$; thus, the latter is not particularly pathological, consistent with our discussion in Section 4.2.

### v3

@@ 4.3 Bias-aware inference
where the bias-aware critical value $\operatorname{cv}_{1-a}(b)$ is given by the number such that $r(b;\operatorname{cv}_{1-a}(b))=a$, and $r(\cdot,\cdot)$ is defined in Corollary˜3.1. By construction, this bias-aware confidence interval has correct (but potentially conservative) asymptotic coverage.
@@
Impose Assumption˜2.1, $\zeta=1/2$, and $\operatorname{aVar}(\hat{\delta}_{h})>0$. Then
@@
The result follows immediately from Propositions˜3.2 and 4.1. ∎
It turns out, however, that a very tight bound $M$ on the signal-to-noise ratio is required for the bias-aware VAR interval to be shorter than the LP interval. Figure˜4.3 plots the relative interval length as a function of the relative asymptotic standard deviation of VAR and LP, for a significance level of $a=10\%$ and for different misspecification bounds $M$. The figure shows that $M$ has to be quite small—apparently below 1—for the bias-aware VAR length to dominate the LP length regardless of the DGP and horizon. Even for $M=1.5$, bias-aware VAR is at best only moderately shorter than LP. Finally, for values of $M$ above 2 (corresponding to a noise-to-signal ratio above $4/T$), bias-aware VAR is dominated by LP.
Figure 4.3: Relative length of bias-aware VAR confidence interval vs. conventional LP interval. Significance level $a=10\%$. Horizontal axis: relative asymptotic standard deviation of VAR vs. LP. Different lines: different bounds $M$ on $\|\alpha(L)\|$. The solid horizontal line marks the value 1.
In Section˜A.2 we furthermore show that the conventional LP confidence interval is at worst slightly wider than a more efficient bias-aware confidence interval centered at the model averaging estimator $\hat{\theta}_{h}(\omega)=\omega\hat{\beta}_{h}+(1-\omega)\hat{\delta}_{h}$, introduced in Corollary˜4.2 above. Even if the weight $\omega$ is chosen to optimize confidence interval length, the gains relative to the LP interval are very small when $M\geq 2$ (corresponding to a noise-to-signal ratio above $4/T$).
@@ 4.4 Inference on multiple impulse responses
Since the least favorable MA polynomial derived in Section˜4.2 depends on the horizon $h$ of interest, one might hope that VARs would not be as prone to bias and thereby undercoverage if interest centers on multiple impulse responses. Unfortunately, LABEL:app:joint shows that this is not the case. There we consider inference on a vector of impulse responses for any combination of response variables $i$, shocks $j$, and horizons $h$. Generalizing Proposition˜4.1, we show that the worst-case norm of the bias is non-negligible if the VAR offers efficiency gains for any linear combination of the parameters of interest. This implies that the conventional VAR confidence interval has fragile coverage even if the target parameter is a linear combination of impulse responses (such as the integral or sum across multiple horizons, as in the fiscal multiplier applications reviewed in Ramey, 2016). The conventional Wald confidence ellipsoid centered at the VAR estimator is similarly fragile.
5 Practical relevance
This section establishes the practical relevance of our theoretical conclusions by comprehensively reviewing current practice for lag length selection in the applied VAR literature, coupled with a simulation study calibrated to the application in Känzig (2021).
5.1 Review of current practice for VAR lag length selection
To evaluate lag length selection practice in the applied VAR literature, we created a comprehensive list of articles published between January 2015 and June 2025 in six top economics journals: American Economic Review, American Economic Journal: Macroeconomics, Econometrica, Journal of Political Economy, Quarterly Journal of Economics, and Review of Economic Studies. We restrict attention to papers that use VARs on time series data (not panel data) to estimate structural impulse response functions. This yields 81 papers total. For each paper, we picked a single specification as the “main” one. [^9] In a few cases, the main use of VARs in a paper was only reported in the appendix. To be conservative, if multiple specifications received equal attention in the main analysis, we picked the one with the largest lag length. The full list of papers, together with the recorded information about the VAR specifications, is provided in the online replication materials.
Our findings suggest that the theoretical results of the preceding sections are likely to have bite in practice, as typical VAR estimation lag lengths in applied papers are short or moderate. The modal lag length is 4 in quarterly data and 12 in monthly data, the mean is slightly below the mode, and less than 10% of papers employ lag lengths greater than or equal to twice the modal values. [^10] Additionally, all the various empirical VAR specifications reported in the handbook chapters of Ramey (2016) and Stock and Watson (2016) are consistent with these patterns. 20% of papers select the lag length in a data-dependent way, often using information criteria. On average across papers, the estimation lag length is only 28% as large as the the longest reported impulse horizon. We conclude that few applied papers follow the recommendation of Kilian and Lütkepohl (2017, pp. 58–66) to use long lag lengths and avoid information criteria. This suggests that VAR inference results reported in much of the applied literature could be subject to the fragility we highlighted in the preceding sections. In fact, since around 40% of papers employ Bayesian shrinkage, the estimation bias could be even larger than indicated by the lag length alone.
5.2 Empirically calibrated simulation study
We now show through simulations that our asymptotic results are informative about the performance of LPs and VARs in an empirically relevant finite-sample setting when the lag length is selected as in current applied practice. Our DGP is calibrated to the oil news shock application in Känzig (2021).
@@ 5.2 Empirically calibrated simulation study
The DGP is a VAR estimated on the dataset of Känzig (2021), using a somewhat longer lag length than that used in the paper. The data series are that paper’s oil shock proxy, the real price of oil, world oil production, world oil inventories, world industrial production, U.S. industrial production, and the U.S. consumer price index (CPI). Whereas Känzig employs 12 lags, we estimate a recursively identified VAR($18$) by OLS on his data and use this as the simulation DGP, with i.i.d. Gaussian shocks. We do not claim that the VAR(18) DGP is more “realistic” than a VAR(12) estimated on the same data, but we contend that it is desirable that confidence intervals should have reliable coverage in both these DGPs.
The parameters of interest are the impulse responses of CPI to the observed oil shock. To be consistent with our theory, we use an “internal instruments” specification that orders the proxy first in the VAR; this differs from the “external instruments” specification used by Känzig. The sample size is $T=720$ months. We will entertain different choices of the estimation lag length $p$ for the LP and VAR estimators. We report results for both delta method and bootstrap confidence intervals. Results are based on 10,000 Monte Carlo simulations. See LABEL:app:oil for implementation details.
@@ Results.
Figure˜5.1 shows that our theoretical results on LP robustness and VAR fragility are practically relevant. The figure depicts the coverage probabilities (left panel) and median confidence interval length (right panel) for VARs (in red, solid and dashed) and LPs (in blue, solid and dashed). The top panel fixes the estimation lag length at $p=12$, while the bottom panel selects the lag length by AIC. Both these choices are frequently encountered in the applied literature, as documented earlier. Given these conventional lag lengths, VAR confidence intervals tend to be shorter than LP intervals, but quite materially undercover, with coverage falling below 60% at medium and long horizons. LP instead attains close to the nominal coverage level of 90% throughout, as expected. The mean lag length selected by AIC is $9.7$, evidently insufficient to guard against dynamic misspecification.
Lag length $p=12$
@@
Figure 5.1: Coverage probability (left) and median length (right) for VAR (red) and LP (blue) nominal 90% confidence intervals computed via the delta method or bootstrap (the latter are indicated with subscript “b” in the legends). Lag length: fixed at $p=12$ in the top panel, and selected using AIC in the bottom panel.
LABEL:app:oil illustrates that the VAR under-coverage can be ameliorated by increasing the lag length beyond what is typically used in current applied practice, at the expense of higher variance, consistent with Section˜3.3. The supplement also reports that the VAR estimator achieves lower MSE than LP, suggesting that—despite the poor performance of the VAR confidence interval—the larger bias of the VAR estimator may not compromise its usefulness as a point estimator (see also Li, Plagborg-Møller, and Wolf, 2024).
Montiel Olea, Plagborg-Møller, Qian, and Wolf (2025) find that the qualitative conclusions above extend to a wide range of empirically calibrated simulation DGPs based on richly specified dynamic factor models: VAR confidence intervals fail to adequately control coverage in a sizable fraction of DGPs, while LP confidence intervals robustly maintain accurate coverage. They document that LP is not only more robust to lag length selection, but also to the choice of control variables, consistent with the theory in this paper.